OpenStudy (anonymous):

solve each system x+2/6 - y+4/3+z/2=0 x+1/2+y-1/2-z/4=9/2 x-5/4+y+1/3+z-2/2=19/4

OpenStudy (anonymous):

I was taught to use the elimination method with these ones. \[\frac{ x+2 }{ 6 }-\frac{ y+4 }{ 3 }+\frac{ z }{ 2 }=0\] Get rid of the fractions by multipling the denominator by it's HCF (Highest Common Factor).

OpenStudy (anonymous):

Do that for all of the three equations. Tell me what you get when you did that.

OpenStudy (anonymous):

So for the first one 6 is highest so I would get X+2+2(y+4)+3z=0 ?

OpenStudy (anonymous):

-2(y+4)*

OpenStudy (anonymous):

remember that minus sign. You don't wave a magic wand and make it disappear.

OpenStudy (anonymous):

Right!!

OpenStudy (anonymous):

But that's good.

OpenStudy (anonymous):

So what would you get for the second one.

OpenStudy (anonymous):

?*

OpenStudy (anonymous):

4 would be highest so 2(x+1)+2(y-1)-z=18 ?

OpenStudy (anonymous):

How come you got a minus sign for z this time?

OpenStudy (anonymous):

Wrote second one wrong should be x+1/2+y-1/2-z/4=9/2

OpenStudy (anonymous):

Ah okay. Then well done correct.

OpenStudy (anonymous):

Do the same thing for the third equation.

OpenStudy (anonymous):

After that just expand all three equations and collect the like terms. Tell me what you get for each equation.

OpenStudy (anonymous):

? Lol kinda lost now

OpenStudy (anonymous):

Show me the equatons that had no fractions.

OpenStudy (anonymous):

equations*

OpenStudy (anonymous):

\[x+2-2(y+4)+3z=0\] \[x+2-2y-8+3z=0\] \[x-2y+3z=-2+8\] \[x-2y+3z=6[1]\] There's your first equation. I placed a number inside some brackets to indicate that this is your first equation. (ie. [1])

OpenStudy (anonymous):

Are you able to put the other two equations in that format please? Show me what you get after doing that.

OpenStudy (anonymous):

@ninab731

OpenStudy (anonymous):

2x+2y-z=17 ?

OpenStudy (anonymous):

Close but not quite.

OpenStudy (anonymous):

Expand all the brackets \[2x+2+2y-2-z=18\] \[2x+2y-z=18-2+2\] \[2x+2y-z=18[2]\] Second equation done.

OpenStudy (anonymous):

Try and do the third equation.

OpenStudy (anonymous):

3x+4y+6z=47

OpenStudy (anonymous):

Oops I goofed

OpenStudy (anonymous):

Not quite.

OpenStudy (anonymous):

Multiply by 12 and expand. \[3x-15+4y+4+6z-12=57\] \[3x+4y+6z=57+15-4+12\] \[3x+4y+6z=34[3]\]

OpenStudy (anonymous):

Third equation done.

OpenStudy (anonymous):

Now Let me get all three equations listed here. \[x-2y+3z=6[1]\] \[2x+2y-z=18[2]\] \[3x+4y+6z=32[3]\]

OpenStudy (anonymous):

What you want is to get rid of one variable and then after that, you try and get rid of the last variable.

OpenStudy (anonymous):

The easiest variable to get rid of is the y. Because all you hav e to do is multipl by 2 in the first 2 equations and then once you've done that, you can subtract each equation by another equation in order to get rid of the y.

OpenStudy (anonymous):

So my first step would be multiplying the first equation by 2. It's the same equation but the numbers are just twice as big. \[[1]\times 2\] \[2(x-2y+3z)=2(6)\]

OpenStudy (anonymous):

Do you get me so far?

OpenStudy (anonymous):

Got it

OpenStudy (anonymous):

Okay now we expand. Even though this equation is practically the same as the first, we will rename it and say that this is the fourth equation [4]. \[2x-4y+3z=12[4]\]

OpenStudy (anonymous):

Okay, I multiplied by 2 so then I turn the -2y into -4y. The 4 in the -4y is what we want because there's a 4y in equation [3].

OpenStudy (anonymous):

SO if we add equation [4] with equation [3] we can ELIMINATE the 4y and -4y so then there will only be two variables. You get me so far?

OpenStudy (anonymous):

Let me show you.

OpenStudy (anonymous):

\[[4]+[3]\] \[(2x−4y+3z)+(3x+4y+6z)=(12)+(32)\] We collect like terms and then I will call this equation [5]. \[5x+9z=44[5]\]

OpenStudy (anonymous):

Do you see what I did there?

OpenStudy (anonymous):

@ninab731

OpenStudy (anonymous):

Ya

OpenStudy (anonymous):

Okay. Try and do the same for equation 2.

OpenStudy (anonymous):

Multiply both sides by 2 to get a 4y. and then subtract the equation with equation 3.

OpenStudy (anonymous):

@ninab731

OpenStudy (anonymous):

If then I multiply by -2 I get -4x-4y+2z=-36

OpenStudy (anonymous):

Don't multiply by a negative. I told you multiply by 2.

OpenStudy (anonymous):

4x+4y-2z=36

OpenStudy (anonymous):

Okay. Equation three is this right? And we want to get rid of the 4y. \[3x+4y+6z=32[3]\] Would be subtract equation two with equation three or would be add them both together?

OpenStudy (anonymous):

Would we*

OpenStudy (anonymous):

Subtract

OpenStudy (anonymous):

Good!! Excellent.

OpenStudy (anonymous):

X-8z=2 ?

OpenStudy (anonymous):

Well done.

OpenStudy (anonymous):

Perfect!

OpenStudy (anonymous):

Now we have two equations with two variables. \[5x+9z=44[5]\] I just made this equation [6]. \[x-8z=2[6]\]

OpenStudy (anonymous):

Now we want to get rid of another variable.

OpenStudy (anonymous):

We can use substitution because we can already make x the subject in equation 6 without dividing or anything like that.

OpenStudy (anonymous):

So \[x=8z+2\] Right?

OpenStudy (anonymous):

Ah ok

OpenStudy (anonymous):

sub it into equation 5.

OpenStudy (anonymous):

and you can find z

OpenStudy (anonymous):

And then once you do that. sub that z-value into either equation 5 or 6.

OpenStudy (anonymous):

40z+10+9z=44 49z+10=44

OpenStudy (anonymous):

Yep good job.

OpenStudy (anonymous):

You can find x by putting what you get for z into one of the two equations that have two variables (ie. [5] or [6])

OpenStudy (anonymous):

But then 44-10=34 and when u divide 49 get decimal

OpenStudy (anonymous):

Do you have the correct answers that say differently?

OpenStudy (anonymous):

You sub in the fraction.

OpenStudy (anonymous):

Don't round up as a decimal. use the fraction.

OpenStudy (anonymous):

Suppose to be (4,8,6)

OpenStudy (anonymous):

rewrite all three equations. because you said equation two had a minus. Can you rewrite the three equations you were given because you must of done something.

OpenStudy (anonymous):

Equations are x+2/6 - y+4/3 + z/2 =0 x+1/2 + y-1/2 - z/4 =9/2 x-5/4 + y+1/3 + z-2/2 =19/4

OpenStudy (anonymous):

\[4x+4y-2z=36[1]\] \[2x-4y+6z=12[2]\] \[3x+4y+6z=80[3]\] \[[1]-[3]\] \[(4x+4y-2z)-(3x+4y+6z)=(36)-(80)\] \[x-8z=-44[4]\] \[[2]+[3]\] \[(2x-4y+6z)+(3x+4y+6z)=(12)+(80) \[5x+12z=92[5]\] \[x=8z-44\] \[5(8z-44)+12z=92\] \[40z-220+12z=92\] \[52z=312\] \[z=6\]

OpenStudy (anonymous):

if three are unknowns n two given syatem then syatem has no solution